# Structured and Balanced Multi-component and Multi-layer Networks

### v1

@@ Abstract
In this work, we propose a balanced multi-component and multi-layer neural network (MMNN) structure to approximate functions with complex features with both accuracy and efficiency in terms of degrees of freedom and computation cost. The main idea is motivated by a multi-component, each of which can be approximated effectively by a single-layer network, and multi-layer decomposition in a “divide-and-conquer” type of strategy to deal with a complex function. While an easy modification to fully connected neural networks (FCNNs) or multi-layer perceptrons (MLPs) through the introduction of balanced multi-component structures in the network, MMNNs achieve a significant reduction of training parameters, a much more efficient training process, and a much improved accuracy compared to FCNNs or MLPs. Extensive numerical experiments are presented to illustrate the effectiveness of MMNNs in approximating high oscillatory functions and its automatic adaptivity in capturing localized features. Our codes and implementation details are available here.
Key words. deep neural networks, rectified linear unit, function compositions, Fourier series
@@ 2 Multi-component and multi-layer neural network (MMNN)
This section begins with an overview of the main notations used in this paper, as detailed in Section 2.1. Subsequently, we introduce a novel network architecture, the Multi-component and Multi-layer Neural Network (MMNN), which is developed based on the balanced decomposition principle discussed in Section 2.2. Following this, in Section 2.3, we outline the learning strategy of MMNN and highlight its advantages over other methods. Finally, in Section 2.4, we compare the numerical performance of MMNNs and FCNNs.
2.1 Notations
The following is an overview of the basic notations used in this paper.
  • The symbols ${\mathbb{N}}$, ${\mathbb{Z}}$, ${\mathbb{Q}}$, and ${\mathbb{R}}$ are used to denote the sets of natural numbers (including 00), integers, rational numbers, and real numbers, respectively. The set of positive natural numbers is denoted as +=\$\0{\mathbb{N}}^{+}={\mathbb{N}}\backslash\{0\}$.
  • The indicator (or characteristic) function of a set $A$, denoted by ${\mathds{1}}_{A}$, is a function that takes the value $1$ for elements of $A$ and 00 for elements not in $A$.
  • The floor and ceiling functions of a real number $x$ can be represented as $\lfloor x\rfloor=\max\{n:n\leq x,\ n\in{\mathbb{Z}}\}$ and $\lceil x\rceil=\min\{n:n\geq x,\ n\in{\mathbb{Z}}\}$.
  • Vectors are denoted by bold lowercase letters, such as ${\bm{a}}=(a_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},a_{n})\in{\mathbb{R}}^{n}$. On the other hand, matrices are represented by bold uppercase letters. For example, $\bm{A}\in\mathbb{R}^{m\times n}$ refers to a real matrix of size $m\times n$, and $\bm{A}^{\textsf{T}}$ denotes the transpose of matrix $\bm{A}$.
  • Slicing notation is used for a vector ${\bm{x}}=(x_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},x_{d})\in{\mathbb{R}}^{d}$, where ${\bm{x}}{[n:m]}$ denotes a slice of ${\bm{x}}$ from its $n$-th to the $m$-th entries for any $n,m\in\{1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},d\}$ with $n\leq m$ and ${\bm{x}}{[n]}$ denotes the $n$-th entry of ${\bm{x}}$. For example, if ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathbb{R}}^{3}$, then $(5{\bm{x}}){[2:3]}=(5x_{2},5x_{3})$ and $(6{\bm{x}}+1){[3]}=6x_{3}+1$. A similar notation is employed for matrices. For instance, ${\bm{A}}[:,i]$ refers to the $i$-th column of ${\bm{A}}$, whereas ${\bm{A}}[i,:]$ indicates the $i$-th row of ${\bm{A}}$. Additionally, ${\bm{A}}[i,n:m]$ corresponds to $({\bm{A}}[i,:])[n:m]$, which means it extracts the entries from the $n$-th to the $m$-th in the $i$-th row.
@@ 2.2 Architecture of MMNNs
In this section, we introduce our Multi-component and Multi-layer Neural Network (MMNN). Each layer of MMNN is a (shallow) neural network of the form
@@
to approximate a vector-valued function ${\bm{f}}\in C([0,1]^{d_{\textnormal{in}}};{\mathbb{R}}^{d_{\textnormal{out}}})$, where ${\bm{W}}\in{\mathbb{R}}^{n\times{d_{\textnormal{in}}}},{\bm{A}}\in{\mathbb{R}}^{{d_{\textnormal{out}}}\times n}$, and $n$ is the width of this network. Here, $\sigma:\mathbb{R}\to\mathbb{R}$ represents the activation function that can be applied elementwise to vector inputs. Throughout this paper, the activation function is $\mathtt{ReLU}$, unless otherwise specified. One can also write it in a more compact form,
${\bm{h}}={\bm{A}}\sigma({\bm{W}}{\bm{x}}+{\bm{b}})+{\bm{c}}={\widetilde{\bm{A}}}\left[\begin{matrix}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})\\ 1\end{matrix}\right]={\widetilde{\bm{A}}}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}}),$ (1)
@@
We call each element of ${\bm{h}}$, i.e., ${\bm{h}}{[i]}=\widetilde{\bm{A}}{[i,:]}\cdot\left[\begin{matrix}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})\\ 1\end{matrix}\right]$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},d_{out}$, a component. Here are a few key features of ${\bm{h}}$:
@@
In comparison, each layer in a typical deep FCNN takes the form $\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$, and each hidden neuron is individually a function of the input ${\bm{x}}$ or each point ${\bm{x}}\in{\mathbb{R}}^{d_{in}}$ is mapped to ${\mathbb{R}}^{n}$, where $n$ is the layer width. All weights ${\widetilde{\bm{W}}}$ are training parameters. In MMNN, each layer is composed of multiple components ${\widetilde{\bm{A}}}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$. Each component is a linear combination of randomly parameterized hidden neurons $\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$, which can be more effectively and stably trained through ${\widetilde{\bm{A}}}$ as a smooth decomposition/transformation. Typically the number of components $d_{out}$ is (much) smaller than the layer width $n$ in our experiments.
A MMNN is a multi-layer composition of ${\bm{h}}_{i}$, i.e., ${\bm{h}}:{\mathbb{R}}^{d_{\textnormal{in}}}\mapsto{\mathbb{R}}^{d_{\textnormal{out}}}$
@@
The width of this MMNN is defined as $\max\{n_{i}:i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},m-1\}$, the rank as $\max\{d_{i}:i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},m-1\}$, and the depth as $m$. To simplify, we denote a network with width $w$, rank $r$, and depth $l$ using the compact notation $(w,r,l)$. See Figure 1(a) for an illustration of MMNN of size $(4,2,2)$. In contrast, an FCNN ${\bm{\phi}}$ can be expressed in the following composition form
@@
The above definition of ResMMNNs requires $d_{1}=\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}=d_{m-1}=r$. If this condition does not hold, we can alternatively define ResMMNN via
@@
${\bm{f}}\oplus{\bm{g}}\coloneqq({\widetilde{\bm{f}}}+{\widetilde{\bm{g}}})[1:d_{\bm{g}}],\quad\textnormal{where}\quad{\widetilde{\bm{f}}}=\begin{bmatrix}{\bm{f}}\\ {\bm{0}}\end{bmatrix}\in\mathbb{R}^{\max\{d_{\bm{f}},d_{\bm{g}}\}}\quad\textnormal{and}\quad{\widetilde{\bm{g}}}=\begin{bmatrix}{\bm{g}}\\ {\bm{0}}\end{bmatrix}\in\mathbb{R}^{\max\{d_{\bm{f}},d_{\bm{g}}\}}.$
It is noteworthy that such an operation is both straightforward and cost-effective to implement. For example, in Python, one can use the following code:
y = f(x);  z = g(x);  n = min(len(y), len(z));  z[:n] = y[:n] + z[:n]
[…]
@@
(b) FCNN of size $(4,2,3)$, i.e., width $4$ and depth $2$.
@@ 2.3 Learning strategy of MMNNs
Our learning strategy is motivated by the following basic principle: a function can be decomposed in a multi-component and multi-layer structure each component of which can be approximated and trained effectively using a one-hidden-layer network, which is a linear combination of random basis functions (e.g., of the form $\sigma({\bm{W}}_{i}\cdot{\bm{x}}+{\bm{b}}_{i})$, see Section 3). Hence optimizing the linear combination weights of the random basis functions, i.e., ${\bm{A}}_{i},{\bm{c}}_{i}$ is both efficient and adequate. On the other hand, optimizing the weights (orientations of the basis functions) ${\bm{W}}_{i}$’s and biases ${\bm{b}}_{i}$’s to make the basis functions more adaptive to fine features of the target function, which would require capturing high-frequency information by a single layer network, leads to not only significantly more parameters to optimize but also difficulties in training as shown in [38]. Specifically, for each layer of MMNN, we fix the activation function parameters ($\bm{W}_{i}$’s and $\bm{b}_{i}$’s) as per PyTorch’s default setting during the training process. This entails initializing both weights and biases uniformly from the distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$, where $k=\frac{1}{\text{in\_features}}$. [^1] It is noteworthy that this initialization approach is similar to the widely used Xavier initialization [4], which draws weights from the distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$ with $k=\frac{\sqrt{6}}{\text{in\_features}+\text{out\_features}}$ and sets the bias to ${\bm{0}}$. The whole training process optimizes all ${\bm{A}}_{i}$’s and ${\bm{c}}_{i}$’s simultaneously using Adam. Note that it is important to have a uniform sampling of orientations ${\bm{W}}_{i}$ and biases ${\bm{b}}_{i}$ for the random basis functions to be able to approximate an arbitrary smooth function well. Unless stated otherwise, parameter initialization adheres to the default settings provided by PyTorch in our experiments.
@@
MMNN2 (S1) (590, 28, 6) 83331 / 170061 $1.39\times 10^{-5}$ $2.80\times 10^{-2}$ 25.2s / 1000 epochs
@@
2.4 MMNNs versus FCNNs
Previously in Section 1, we discussed the theoretical differences between MMNNs and FCNNs. Now, let’s explore and compare their numerical performance. To ensure a fair comparison, we will use networks with a similar number of parameters, ensuring that all networks have sufficient parameters to learn the target function effectively. Typically, when training an FCNN, all parameters are optimized. For a thorough comparison, we will employ two learning strategies for MMNNs as detailed in Section 2.3: S1 and S2. S1 involves training approximately half the number of parameters of the MMNN, while S2 involves training all parameters.
We choose a 1D function $f_{1}(x)=\cos(20\pi|x|^{1.4})+0.5\cos(12\pi|x|^{1.6})$ and a 2D function
$f_{2}(x_{1},x_{2})=\sum_{i=1}^{2}\sum_{j=1}^{2}a_{ij}\sin(sb_{i}x_{i}+sc_{ij}x_{i}x_{j})\cos(sb_{j}x_{j}+sd_{ij}x_{i}^{2}),$

### v2

# Structured and Balanced Multi-component and Multi-layer Neural Networks

@@ Abstract
In this work, we propose a balanced multi-component and multi-layer neural network (MMNN) structure to accurately and efficiently approximate functions with complex features, in terms of both degrees of freedom and computational cost. The main idea is inspired by a multi-component approach, in which each component can be effectively approximated by a single-layer network, combined with a multi-layer decomposition strategy to capture the complexity of the target function. Although MMNNs can be viewed as a simple modification of fully connected neural networks (FCNNs) or multi-layer perceptrons (MLPs) by introducing balanced multi-component structures, they achieve a significant reduction in training parameters, a much more efficient training process, and improved accuracy compared to FCNNs or MLPs. Extensive numerical experiments demonstrate the effectiveness of MMNNs in approximating highly oscillatory functions and their ability to automatically adapt to localized features. Our code and implementations are available at GitHub.
keywords:
structured decomposition, function compositions, deep neural networks, rectified linear unit, Fourier analysis
@@ 2 Multi-component and multi-layer neural network (MMNN)
In this section, we present a novel network architecture called the Multi-component and Multi-layer Neural Network (MMNN). Let’s begin with some notations. Let ${\mathbb{R}}$ represent the set of real numbers. The indicator (or characteristic) function of a set $A$, denoted by ${\mathds{1}}_{A}$, is a function that takes the value $1$ for elements in $A$ and 00 for elements not in $A$. Vectors and matrices are denoted by bold lowercase and uppercase letters, respectively. We use slicing notation for a vector ${\bm{x}}=(x_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},x_{d})\in{\mathbb{R}}^{d}$, where ${\bm{x}}{[n:m]}$ denotes a slice of ${\bm{x}}$ from its $n$-th to the $m$-th entries for any $n,m\in\{1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},d\}$ with $n\leq m$, and ${\bm{x}}{[n]}$ denotes the $n$-th entry of ${\bm{x}}$. For example, if ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathbb{R}}^{3}$, then $(5{\bm{x}}){[2:3]}=(5x_{2},5x_{3})$ and $(6{\bm{x}}+1){[3]}=6x_{3}+1$. A similar notation is used for matrices. For instance, ${\bm{A}}[:,i]$ refers to the $i$-th column of ${\bm{A}}$, whereas ${\bm{A}}[i,:]$ indicates the $i$-th row of ${\bm{A}}$. Additionally, ${\bm{A}}[i,n:m]$ corresponds to $({\bm{A}}[i,:])[n:m]$, extracting the entries from the $n$-th to the $m$-th in the $i$-th row.
Later in this section, we introduce the architecture of MMNNs in Section 2.1. Following this, in Section 2.2, we outline the learning strategy of MMNN and highlight its advantages over other methods. Finally, in Section 3, we compare the numerical performance of MMNNs and FCNNs.
@@ 2.1 Architecture of MMNNs
In this section, we introduce the architecture of our Multi-component and Multi-layer Neural Network (MMNN). Each layer of the MMNN is a (shallow) neural network of the form
@@
to approximate a vector-valued function ${\bm{f}}:{\mathbb{R}}^{d_{\textnormal{in}}}\to{\mathbb{R}}^{d_{\textnormal{out}}}$ where ${\bm{W}}\in{\mathbb{R}}^{n\times{d_{\textnormal{in}}}},{\bm{A}}\in{\mathbb{R}}^{{d_{\textnormal{out}}}\times n}$, and $n$ is the width of this network. Here, $\sigma:\mathbb{R}\to\mathbb{R}$ represents the activation function that can be applied elementwise to vector inputs. Throughout this paper, the activation function is $\mathtt{ReLU}$, unless otherwise specified. One can also write it in a more compact form,
(1) ${\bm{h}}={\bm{A}}\sigma({\bm{W}}{\bm{x}}+{\bm{b}})+{\bm{c}}={\widetilde{\bm{A}}}\left[\begin{matrix}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})\\ 1\end{matrix}\right],$
@@
We call each element of ${\bm{h}}$, i.e., ${\bm{h}}{[i]}=\widetilde{\bm{A}}{[i,:]}\cdot\left[\begin{matrix}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})\\ 1\end{matrix}\right]$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},d_{\textnormal{out}}$, a component. Here are a few key features of ${\bm{h}}$:
@@
In comparison, each layer in a typical deep FCNN takes the form $\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$, and each hidden neuron is individually a function of the input ${\bm{x}}$ or each point ${\bm{x}}\in{\mathbb{R}}^{d_{\textnormal{in}}}$ is mapped to ${\mathbb{R}}^{n}$, where $n$ is the layer width. All weights ${\widetilde{\bm{W}}}$ are training parameters. In MMNNs, each layer is composed of multiple components ${\widetilde{\bm{A}}}\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$. Each component is a linear combination of randomly parameterized hidden neurons $\sigma({\widetilde{\bm{W}}}{\widetilde{\bm{x}}})$, which can be more effectively and stably trained through ${\widetilde{\bm{A}}}$ as a smooth decomposition/transformation. Typically the number of components $d_{\textnormal{out}}$, the dimension of intermediate feature space, is (much) smaller than the layer width $n$, the number of random neurons (or basis functions) in the MMNN. On one hand, the intermediate feature space is compressed, and on the other hand, there are diverse random basis whose linear combinations can have enough representation power in the feature space.
An MMNN is a multi-layer composition of ${\bm{h}}_{i}$, i.e., ${\bm{h}}:{\mathbb{R}}^{d_{\textnormal{in}}}\mapsto{\mathbb{R}}^{d_{\textnormal{out}}}$
@@
The width of this MMNN is defined as $\max\{n_{i}:i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},m-1\}$, the rank as $\max\{d_{i}:i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},m-1\}$, and the depth as $m$. MMNNs are designed to reduce the dimensionality of the intermediate feature space, thereby simplifying optimization while largely preserving the network’s expressive power. To simplify, we denote a network with width $w$, rank $r$, and depth $l$ using the compact notation $(w,r,l)$. See Figure 1(a) for an illustration of an MMNN of size $(4,2,2)$. In contrast, an FCNN ${\bm{\phi}}$ can be expressed in the following composition form
@@
We clarify the structural differences between FCNNs and MMNNs, omitting bias terms for notational simplicity. In MMNNs, each layer is defined by a composition
${\bm{h}}:\mathbb{R}^{r}\to\mathbb{R}^{r},\quad{\bm{h}}({\bm{x}})={\bm{A}}\sigma({\bm{W}}{\bm{x}}),$
while in standard FCNNs, layers are typically written as
${\bm{h}}:\mathbb{R}^{n}\to\mathbb{R}^{n},\quad{\bm{h}}({\bm{x}})=\sigma({\bm{W}}{\bm{x}}),$
where $r\ll n$, and $r$ denotes the MMNN rank, an internal dimensionality that helps regulate network complexity. The core principle behind MMNNs is to limit the dimensionality of intermediate space, which helps streamline the optimization process while maintaining the model’s expressive power.
One might suggest that the product ${\bm{A}}_{j}{\bm{W}}_{j+1}$ in MMNNs can be collapsed into a single matrix, making MMNNs a special case of FCNNs. However, this overlooks two critical distinctions:
  1. (1)
    Asymmetric Parameter Roles: In MMNNs, ${\bm{W}}$ is randomly initialized and fixed while ${\bm{A}}$ is learnable. This asymmetry is central to our representation and training strategy and cannot be replicated by simply reinterpreting MMNNs as FCNNs with a low rank factorization of the weight matrix ${\bm{W}}$ which needs to be learned fully posing a challenging task for the optimization.
  2. (2)
    Architectural Interventions: In practice, modern networks often employ techniques such as batch normalization, dropout, and residual connections, which fundamentally alter the layer-wise composition. For instance,
    • With batch normalization layers $B_{i}$’s:
      ${\bm{h}}_{m}\circ B_{m-1}\circ{\bm{h}}_{m-1}\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\circ B_{1}\circ{\bm{h}}_{1};$
    • With dropout layers $D_{i}$’s:
      ${\bm{h}}_{m}\circ D_{m-1}\circ{\bm{h}}_{m-1}\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\circ D_{1}\circ{\bm{h}}_{1};$
    • With residual connections:
      ${\bm{h}}_{m}\circ({\bm{I}}+{\bm{h}}_{m-1})\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\circ({\bm{I}}+{\bm{h}}_{2})\circ{\bm{h}}_{1}.$
    In these scenarios, the clean compositional structure required to merge adjacent matrices (e.g., ${\bm{A}}_{j}$ and ${\bm{W}}_{j+1}$) is disrupted. Thus, MMNNs remain distinct in both form and training philosophy.
@@
The above definition of ResMMNNs requires $d_{1}=\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}=d_{m-1}=r$. If this condition does not hold, we can alternatively define ResMMNNs via
@@
${\bm{f}}\oplus{\bm{g}}\coloneqq({\widetilde{\bm{f}}}+{\widetilde{\bm{g}}})[1:d_{\bm{g}}],$
where
${\widetilde{\bm{f}}}=\begin{bmatrix}{\bm{f}}\\ {\bm{0}}\end{bmatrix}\in\mathbb{R}^{\max\{d_{\bm{f}},d_{\bm{g}}\}}\quad\textnormal{and}\quad{\widetilde{\bm{g}}}=\begin{bmatrix}{\bm{g}}\\ {\bm{0}}\end{bmatrix}\in\mathbb{R}^{\max\{d_{\bm{f}},d_{\bm{g}}\}}.$
@@
(b) FCNN of size (4, –, 2), i.e., width $4$ and depth $2$.
@@ 2.2 Learning strategy of MMNNs
Our learning strategy is motivated by the following basic principle: a function can be decomposed in a multi-component and multi-layer structure each component of which can be approximated and trained effectively using a one-hidden-layer network, which is a linear combination of random basis functions (e.g., of the form $\sigma({\bm{W}}_{i}\cdot{\bm{x}}+{\bm{b}}_{i})$, see Section 4). Therefore, optimizing the linear combination weights of the random basis functions, namely ${\bm{A}}_{i}$’s and ${\bm{c}}_{i}$’s, is both computationally efficient and sufficiently expressive. On the other hand, optimizing the weights (orientations of the basis functions) ${\bm{W}}_{i}$’s and biases ${\bm{b}}_{i}$’s to make the basis functions more adaptive to fine-tune features of the target function, which would require capturing high-frequency information by a single layer network, leads to not only significantly more parameters to optimize but also difficulties in training as shown in [41]. Specifically, for each layer of an MMNN, we fix the activation function parameters ($\bm{W}_{i}$’s and $\bm{b}_{i}$’s) as per PyTorch’s default setting during the training process. This entails initializing both weights and biases uniformly from the distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$, where $k=\frac{1}{\text{in\_features}}$. [^1] It is noteworthy that this initialization approach is similar to the widely used Xavier initialization [5], which draws weights from the distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$ with $k=\frac{\sqrt{6}}{\text{in\_features}+\text{out\_features}}$ and sets the bias to ${\bm{0}}$. The whole training process optimizes all ${\bm{A}}_{i}$’s and ${\bm{c}}_{i}$’s simultaneously using the Adam optimizer [16]. Note that it is important to have a uniform sampling of orientations ${\bm{W}}_{i}$ and biases ${\bm{b}}_{i}$ for the random basis functions to be able to approximate an arbitrary smooth function well. Unless stated otherwise, parameter initialization adheres to the default settings provided by PyTorch in our experiments.
@@
MMNN2 (S1) (590, 28, 6) 83331 / 170061 $\bm{1.39\times 10^{-5}}$ $\bm{2.80\times 10^{-2}}$ 25.2s / 1000 epochs
@@
3 MMNNs versus FCNNs
In Section 2.1, we outlined the distinctions between MMNNs and FCNNs regarding their representation and learning approaches. Here, we evaluate their numerical performance for approximating oscillatory functions in Section 3.1 and solving PDEs in Section 3.2. To ensure a fair comparison, we use networks with a similar number of parameters and ensure that all networks have sufficient parameters to learn the target function. Typically, when training an FCNN, all parameters are optimized. For a thorough comparison, we will employ two learning strategies for MMNNs as detailed in Section 2.2: S1 and S2. S1 involves training approximately half the number of parameters of the MMNN, while S2 involves training all parameters.
3.1 Oscillatory function approximation
We choose a one-dimensional function $f_{1}(x)=\cos(20\pi|x|^{1.4})+0.5\cos(12\pi|x|^{1.6})$ and a two-dimensional function
$f_{2}(x_{1},x_{2})=\sum_{i=1}^{2}\sum_{j=1}^{2}a_{ij}\sin(sb_{i}x_{i}+sc_{i,j}x_{i}x_{j})\cos(sb_{j}x_{j}+sd_{i,j}x_{i}^{2}),$